Tohban Report 2015-01-28
From RHESSI Wiki
|Start Date:||21 Jan 2015|
|End Date:||28 Jan 2015|
|Next Tohban:||Nicole Duncan|
|List all reports|
Solar activity is moderate to high and is picking up. There are 9 active regions on the disk, with 2-3 of them flaring. Many of these are coming from AR 12268 at disk center, including an M class flare. There is also a prolific AR rotating onto the disk; this AR produced many flares on its last turn on the merry-go-round and is already showing activity this time around.
How many GOES flares occurred?
Flares above B, C, M, X class were 11 37 2 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 5 25 1 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 0 0 0 0
There were RHESSI flares/GOES flares 277 / 50 over the time range 21-Jan-15 28-Jan-15
We've recently switched to Active / Vigorous decimation, at 2015-028-18:06 UT this morning. Due to increased activity in the past ~day, the SSR is no longer emptying, and was at 35% this morning (before the pass set).
No gaps this week as of Jan 27 (Tuesday).
- Rear fast counts were high again, so we raised the rear fast threshold by 16 steps, in two sets of 8, on the 2015-01-21 BGS pass starting at ~3:30 PST. The new value is 0x60 (was 0x50). Fast rates as eyeballed from the realtime data were approx 58,000 (pre-adjustment), 12,000 (after 8 steps up), and 1600 (after 16 steps up). Livetime percentages were approx 60%, 90%, and 98%, respectively.
15-021-23:33:58 /idpudumptabl table=dibtbl9 15-021-23:34:03 /idputable9 offset=rearfastdac 15-021-23:34:34 /idpuload value=0x58
15-021-23:36:47 /idputable9 offset=rearfastdac 15-021-23:36:55 /idpuload value=0x60
- D6 has been showing some strange peaks in the spectrum; these are best seen as multiple lines (that drift) in spectrograms. On Monday, Jan 28 we raised the slow front threshold on D6 to ~10 keV and kept it there for a day before returning to the nominal ~3 keV. This is intended to see if the slow threshold affects the abnormal spectral lines. For results on this, see the summary that Albert prepared.
15-026-21:57:13 UTC / D6 front slow threshold to 0x30 (was 0x0C). 15-027-21:36:37 UTC / D6 front slow threshold to 0x0C (was 0x30).
- D6 fast rates are a bit high (have been rising slowly over the past week). It would be a good idea to raise the front fast threshold a touch. I didn't want to do this at the same time as the slow threshold experiment.
- A note for the new tohban: the light leak problem continues, and most of the wild departures in the monitor rate plots are from these. The problem is worst in detectors 2, 3, 5, and 9, but also shows up in some of the others. Remember that for D2 this shows up in the fronts because it's unsegmented, while the others usually manifest in the rears.
A spin-up was performed on Monday, Jan 26 from about 14.5 RPM to 14.97 between 20:29 and 22:06 UTC. There were no anomalies.
|Decimation||Active / Vigorous|
|Night time data (fronts)||On|
|Night time data (rears)||On|
|Require extra passes?||No|
|Requirement for moving pointer?||No|